OpenStudy (anonymous):

I seriously need help with question with polynomials. Please help me? Medal and fan!

OpenStudy (anonymous):

A bucket of paint has spilled on a tile floor. The paint flow can be expressed with the function r(t) = 3t, where t represents time in minutes and r represents how far the paint is spreading. The flowing paint is creating a circular pattern on the tile. The area of the pattern can be expressed as A(r) = πr2. Part A: Find the area of the circle of spilled paint as a function of time, or A[r(t)]. Show your work. (6 points) Part B: How large is the area of spilled paint after 10 minutes? You may use 3.14 to approximate π in this problem. (4 points)

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Um, Haseeb? Are you there?

OpenStudy (anonymous):

@Hero If you have the time can you help me out with this?

OpenStudy (beginnersmind):

Can you calculate A for a given value of t? Say t = 2, what will be the painted area?

OpenStudy (anonymous):

Yes.

OpenStudy (beginnersmind):

Ok, so you can solve part B.

OpenStudy (anonymous):

Yes, but Part A is what I'm having trouble with.

OpenStudy (beginnersmind):

Cool. Take the same steps as for part B, but instead of using 10 just keep t. You won't get a number but an expression involving t.

OpenStudy (anonymous):

So it would A[r(t)]=πr2+3t?

OpenStudy (beginnersmind):

Just to check, what result did you get for part B?

OpenStudy (anonymous):

I haven't done Part B yet.

OpenStudy (beginnersmind):

You should. It will help with part A. It also makes it easier to understand someone else's explaination, if you'll still need help.

OpenStudy (anonymous):

Okay, I'll try.

OpenStudy (anonymous):

Wait, but what is the radius of the spilled paint?

OpenStudy (beginnersmind):

r(t) = 3t, where t represents time in minutes and r represents how far the paint is spreading.

OpenStudy (beginnersmind):

so r is meant to be the radius

OpenStudy (anonymous):

Oh, so the paint is spreading equivalent to the radius of the circle?

OpenStudy (beginnersmind):

The radius of the painted circle is increasing as time goes on. So it's 3 after 1 minute, 6 after 2 minutes, etc.

OpenStudy (anonymous):

So it's 282.6?

OpenStudy (beginnersmind):

what did you get for the radius after 10 minutes?

OpenStudy (anonymous):

30.

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

295.788?

OpenStudy (beginnersmind):

Let me check how I can write pretty expressions real quick. :)

OpenStudy (beginnersmind):

\[A =\pi r ^{2}\]

OpenStudy (beginnersmind):

Is that what you used?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

With 30 as r

OpenStudy (anonymous):

Because 30 is the radius after 10 minutes.

OpenStudy (beginnersmind):

When I plug in r = 30 I get\[\pi (30)^{2} = 900\pi= 900x3.14 = 2826\]

OpenStudy (beginnersmind):

Hm, seems like your first answer was almost right, just off by a factor of 10 for some reason.

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

So part B is 2826

OpenStudy (beginnersmind):

Yes. Now for part A.

OpenStudy (beginnersmind):

The idea is the same, but now instead of pluging in 30 into A = πr2 you plug in 3t

OpenStudy (anonymous):

A[r(t)] = pi*3t^2?

OpenStudy (beginnersmind):

More like pi*(3t)^2

OpenStudy (beginnersmind):

Probably should be simplified to 9*pi*t^2

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

And that's part A?

OpenStudy (beginnersmind):

Yes. It gives you an expression to calculate the area directly from time. Without having to calculate the radius first.

OpenStudy (anonymous):

Thank you. :)

OpenStudy (beginnersmind):

No problem :)